# The Application of Tailored Fields for Studying Chirality and Symmetry

### v1

@@ Introduction
Finally, we may contemplate potential novel applications of tailored fields and the direction of this research area as a whole. In our opinion, polarization-tailored sources are especially interesting in the context of non-linear light-matter interactions and should prove highly useful for spectroscopy and coherent control in the coming years. For example, the two-color tailored fields employed to control the polarization of EUV or X-ray light [46, 47, 48, 49, 50, 45, 51, 149, 150, 151, 152] can also be employed to harness valley degrees of freedom [153] and material topology [154]. On the other hand, topological beams can serve as a spectroscopic tool to explore the target’s anisotropy [155] and generate out-of-equilibrium states with novel properties [156], or even control optoelectronic responses [58]. More generally, by taking advantage of the spatially-varying phase in OAM beams, one can hope to control different properties of the optical responses, as has been for instance already employed to control the polarization of the attosecond pulses [157] and their spectral content [158]. Lastly, we highlight one extremely relevant field for applying such capabilities — chirality. Given that chirality connects with the geometrical properties of objects, and tailored fields allow manipulating the geometry of light on multiple levels, the two should match extremely well. In the next section, we introduce chirality and the state-of-art examples of how it benefits from the use of a tailored laser field.
@@
Figure 2: The study of molecular chirality using standard laser fields. Panel A: Scheme illustrating the concept of chirality, a universal property found at all length scales. Molecular enantiomers are the mirror-reflected versions of a chiral molecule. Panels B-D: Traditional methods for the analysis of molecular chirality are based on CD, which measures the difference between the responses of the two enantiomers to elliptically-polarized light. Standard CD consists of measuring differences in absorption of light with right circular polarization (RCP) and left circular polarization (LCP), as illustrated in panel B. This concept can be extended to the x-ray regime, where the enantio-sensitivity is higher, as shown in panel C, figure adapted from 69. CD can also be found in the variation of the HHG yield with the ellipticity of the driving field, as presented in panel D (adapted from 33). Panels E-F: Modern approaches to analyze molecular chirality based on electric-dipole interactions with standard circularly/elliptically polarized light. In PECD, the angular distribution of the photoelectron momentum shows an enantio-sensitive asymmetry in the direction of the laser field propagation, as shown in panel E, reprinted from 159. Enantio-sensitive signals are also found in measurements of photoion fragments (as shown in panel F, reprinted from 160).
@@
Parallel to these findings, in the past years, the emergence of the so-called ”electric-dipole revolution” [166] — i.e. the extraction of huge chiral signals using schemes that rely only on electric-dipole interactions—has launched enormous theoretical and experimental efforts to develop new techniques for chiral sensing, also towards studying gas phase systems and ultrafast chirality. This has been initiated with the emergence of photoelectron CD (PECD), whereby one measures the helicity-dependent angular-, momentum-, and energy-resolved photoemission spectrum. The novelty here is that the chiral signal emerges already within the electric-dipole approximation, reaching scales of up to  30% (orders of magnitude larger than typical CD) [167, 73, 168]. In PECD the mirror-symmetry-broken electric field arises in light’s spatial structure. In that respect, electric-dipole interactions are intuitively expected to not yield chiral signals. However, the experimental apparatus angularly resolves the photoelectron yield, permitting electric-dipole-induced interactions to arise in resolved parts of the spectrum. If one spatially integrates then the chiral signal vanishes, but when looking at the forward/backward hemispheres of the photoemission (with respect to light’s propagation axis) a chiral signal arises in the form of a forward/backward asymmetry (see Figure 2E, reprinted from 159). This effect has also been extended to nonlinear optical interactions in various configurations [169, 170], and even to the attosecond regime [171, 31, 172]. It has also been more recently observed in a similar process of laser-induced electron diffraction [173]. In addition, it has been proposed that photoionization by short few-cycle linearly polarized pulses results in the creation of enantio-sensitive electron vortices [174]. Moreover, a similar concept to PECD consists in Coulomb explosion imaging whereby the photoions are measured in coincidence [175, 160], allowing also to directly measure the absolute handedness of molecules (i.e. if they are in $R/L$ configuration, and not only the presence of dichroism), where symmetry breaking arises due to spatial structure and the detection geometry (see Figure 2F, reprinted from 160).
@@
Figure 3: Applications of tailored laser fields to investigate molecular chirality. Panel A is a figure reprinted from 176 showing a polarization grating creates an enantio-sensitive force that separates molecular enantiomers. Panel B is a figure reprinted from 32, showing an example of how PECD experiments can also benefit from using tailored fields, such as an OTC fields. OTC fields can also be used to drive low-order non-linear interactions where the 3-wave mixing signal is enantiomer-dependent, as shown in panel C, a figure reprinted from 70. Panel D presents a figure reprinted from 177, showing the enantio-specific transfer between molecular rotational states when using a 3-color field with crossed polarizations. Highly non-linear interactions such as HHG can also rise to strong enantio-sensitive signals. Panel E shows a figure reprinted from 79 where the HHG signal induced by the $\omega$$2\omega$ field together with the corresponding CD was presented. Panel F presents a figure reprinted from 166 that illustrates the setup for the generation of locally-chiral light (left side) and shows the enantio-sensitive HHG yield (up right) and corresponding CD (bottom right). Finally, molecular chirality can also be explored using topological light: panel G shows a figure reprinted from 76 exhibiting the HD induced by light carrying OAM, while panel H is a figure adapted from 178 showing the enantio-sensitive spatial distribution of the HHG signal driven by locally-chiral light with OAM.
@@
So, where do tailored light pulses necessarily come in? We would argue that they are pertinent to the field due to the symmetry connection — they permit either breaking symmetries that have not been broken before, or breaking the same symmetries but in a new way. In the following, we shall present a compilation of techniques using tailored light to investigate chirality. One of the first proposals for bringing tailored fields to chirality consisted in the use of polarization gratings that would exert an enatio-sensitive force on chiral molecules, enabling the separation of the enantiomers (see Figure 3A, reprinted from 176). Other complex polarization spatial structures have also demonstrated to result in enantio-sensitive direction of light emission [81, 179].
@@
Let us now move towards all-optical techniques using tailored light. First, low-order non-linear interactions have the advantage of requiring moderate laser intensities. These include second-order optical techniques that break the symmetry in polarization space such as sum-frequency generation or other wave mixing effects [67, 68, 183, 184] (including the microwave regime that has been particularly successful owing to good phase matching [70, 71, 184], as shown in Figure 3C reprinted from 70) or using three resonant pulses that break the mirror symmetry in polarization and time, such as in enantio-specific state transfer measurements (see Figure 3D reprinted from 177) or the enantiosensitive free-induction decay steered by a tricolor cross-polarized field [84].
@@
In the highly non-linear regime, the above-mentioned chiral HHG [33] motivated additional work using tailored light to drive HHG — it was known at the time that combinations of circular drivings allow conditions of nonzero HHG, e.g. if bi-chromatic bi-circular pulses are employed [48]. The hope was to use such poly-chromatic polarization-tailored pulses to repeat this experiment but in conditions where much larger dichroism would be obtained [185]. Indeed, recently dichroism of up to  10% was measured in bicircular HHG [79, 186] (see Figure 3E reprinted from 79), and also utilized to track chiral chemical reactions in real time [187]. In the next stage, perturbative and wave-mixing methods were extended to HHG, where by tailoring the driving beams with polychromatic components the nature of the polarization-induced symmetry breaking can be more finely controlled [78]. Here one can also generate electric-dipole chiral signals that arise in the yield and handedness of the harmonics. Lastly, specialized beams where light’s symmetries are completely broken were developed, denoted as so-called locally-chiral light [53]. The idea here is to impose strict no-mirror symmetry groups on the light field at every local point in space [188, 189]. That is, the time-dependent polarization of the electric (and magnetic) field in every given point in space traces out a spatio-temporal object that cannot be superposed onto its mirror twin. This extends the typical molecular definition of chirality to electromagnetism. A direct result of this polarization-induced symmetry breaking is chiral signals in HHG (and perturbative harmonics as well) that are driven by electric-dipole interactions and arise in the yield of harmonics, rather than in their handedness (see Figure 3F, reprinted from 166). This technique has been proposed in multiple geometries, each with its own advantages/disadvantages [81, 190, 183, 82, 166, 84], and allows all-optical measurements that can be optimized to very high chiral signals, above 100%, and potentially also resolve local stereo-centers within the chiral molecule [191]. Locally-chiral light beams were also proposed as useful for extending PECD methodologies, whereby the geometry of the measurement apparatus is no longer essential to obtain the electric-dipole chiral signal since the light carries the dipole-induced symmetry breaking. This leads to a breaking of the forward/backward asymmetry in PECD and with it chiral signals in the angle-integrated above-threshold ionization spectra that can also be employed for optical enantio-purification [192].
@@
More recently, topological light beams carrying OAM were employed to develop new types of dichroism such as helical dichroism (HD) (see Figure 3G reprinted from 76). This signal appears in the linear optical response as a result of primarily electric quadrupole interactions and is a dichroism obtained upon inverting the beam’s OAM handedness (i.e. the sign of its topological charge) rather than its polarization handedness [74, 75, 76, 77]. Helical dichroism can even appear with linearly polarized beams, decoupling the need for polarization control. Its origin of symmetry breaking is still in light’s spatial structure, but embedded in the phase front rather than polarization evolution. Technically, CD and HD can be combined to provide additional chiral information by controlling both degrees of freedom. In principle, nonlinear responses should also yield to HD (due to the symmetry-breaking nature being the same), which, to our knowledge, has not been generally shown or predicted yet, but only analyzed in nonlinear absorption spectroscopy [193]. Moreover, combining OAM beams with locally-chiral light can produce even more powerful chiral spectroscopy schemes [178] (see Figure 3H reprinted from 178). Similarly, the application of locally-chiral light with evermore intricate waveforms such as three-color pulses could contribute to other chiral sensing methodologies such as enantio-sensitive population transfers [194].

### v2

@@ Introduction
Finally, we may contemplate potential novel applications of tailored fields and the direction of this research area as a whole. In our opinion, polarization-tailored sources are especially interesting in the context of non-linear light-matter interactions and should prove highly useful for spectroscopy and coherent control in the coming years. For example, the two-color tailored fields employed to control the polarization of EUV or X-ray light [46, 47, 48, 49, 50, 45, 51, 149, 150, 151, 152, 153] can also be employed to harness valley degrees of freedom [154, 155] and material topology [156]. On the other hand, topological beams can serve as a spectroscopic tool to explore the target’s anisotropy [157] and generate out-of-equilibrium states with novel properties [158], or even control optoelectronic responses [58]. More generally, by taking advantage of the spatially-varying phase in OAM beams, one can hope to control different properties of the optical responses, as has been for instance already employed to control the polarization of the attosecond pulses [159] and their spectral content [160]. Lastly, we highlight one extremely relevant field for applying such capabilities — chirality. Given that chirality connects with the geometrical properties of objects, and tailored fields allow manipulating the geometry of light on multiple levels, the two should match extremely well. In the next section, we introduce chirality and the state-of-art examples of how it benefits from the use of a tailored laser field.
@@
Figure 2: The study of molecular chirality using standard laser fields. Panel A: Scheme illustrating the concept of chirality, a universal property found at all length scales. Molecular enantiomers are the mirror-reflected versions of a chiral molecule. Panels B-D: Traditional methods for the analysis of molecular chirality are based on CD, which measures the difference between the responses of the two enantiomers to elliptically-polarized light. Standard CD consists of measuring differences in absorption of light with right circular polarization (RCP) and left circular polarization (LCP), as illustrated in panel B. This concept can be extended to the x-ray regime, where the enantio-sensitivity is higher, as shown in panel C, figure adapted from 69. CD can also be found in the variation of the HHG yield with the ellipticity of the driving field, as presented in panel D (adapted from 33). Panels E-F: Modern approaches to analyze molecular chirality based on electric-dipole interactions with standard circularly/elliptically polarized light. In PECD, the angular distribution of the photoelectron momentum shows an enantio-sensitive asymmetry in the direction of the laser field propagation, as shown in panel E, reprinted from 161. Enantio-sensitive signals are also found in measurements of photoion fragments (as shown in panel F, reprinted from 162).
@@
Parallel to these findings, in the past years, the emergence of the so-called ”electric-dipole revolution” [168] — i.e. the extraction of huge chiral signals using schemes that rely only on electric-dipole interactions—has launched enormous theoretical and experimental efforts to develop new techniques for chiral sensing, also towards studying gas phase systems and ultrafast chirality. This has been initiated with the emergence of photoelectron CD (PECD), whereby one measures the helicity-dependent angular-, momentum-, and energy-resolved photoemission spectrum. The novelty here is that the chiral signal emerges already within the electric-dipole approximation, reaching scales of up to  30% (orders of magnitude larger than typical CD) [169, 73, 170]. In PECD the mirror-symmetry-broken electric field arises in light’s spatial structure. In that respect, electric-dipole interactions are intuitively expected to not yield chiral signals. However, the experimental apparatus angularly resolves the photoelectron yield, permitting electric-dipole-induced interactions to arise in resolved parts of the spectrum. If one spatially integrates then the chiral signal vanishes, but when looking at the forward/backward hemispheres of the photoemission (with respect to light’s propagation axis) a chiral signal arises in the form of a forward/backward asymmetry (see Figure 2E, reprinted from 161). This effect has also been extended to nonlinear optical interactions in various configurations [171, 172], and even to the attosecond regime [173, 31, 174]. It has also been more recently observed in a similar process of laser-induced electron diffraction [175]. In addition, it has been proposed that photoionization by short few-cycle linearly polarized pulses results in the creation of enantio-sensitive electron vortices [176]. Moreover, a similar concept to PECD consists in Coulomb explosion imaging whereby the photoions are measured in coincidence [177, 162], allowing also to directly measure the absolute handedness of molecules (i.e. if they are in $R/L$ configuration, and not only the presence of dichroism), where symmetry breaking arises due to spatial structure and the detection geometry (see Figure 2F, reprinted from 162).
@@
Figure 3: Applications of tailored laser fields to investigate molecular chirality. Panel A is a figure reprinted from 178 showing a polarization grating creates an enantio-sensitive force that separates molecular enantiomers. Panel B is a figure reprinted from 32, showing an example of how PECD experiments can also benefit from using tailored fields, such as an OTC fields. OTC fields can also be used to drive low-order non-linear interactions where the 3-wave mixing signal is enantiomer-dependent, as shown in panel C, a figure reprinted from 70. Panel D presents a figure reprinted from 179, showing the enantio-specific transfer between molecular rotational states when using a 3-color field with crossed polarizations. Highly non-linear interactions such as HHG can also rise to strong enantio-sensitive signals. Panel E shows a figure reprinted from 79 where the HHG signal induced by the $\omega$$2\omega$ field together with the corresponding CD was presented. Panel F presents a figure reprinted from 168 that illustrates the setup for the generation of locally-chiral light (left side) and shows the enantio-sensitive HHG yield (up right) and corresponding CD (bottom right). In addition, molecular chirality can also be explored using topological light: panel G shows a figure reprinted from 76 exhibiting the HD induced by light carrying OAM. Finally, locally-chiral light also serves as a tool to imprint chirality on atoms, as shown in H, which is a figure adapted from 85 presenting the chiral superposition of atomic electronic states.
@@
So, where do tailored light pulses necessarily come in? We would argue that they are pertinent to the field due to the symmetry connection — they permit either breaking symmetries that have not been broken before, or breaking the same symmetries but in a new way. In the following, we shall present a compilation of techniques using tailored light to investigate chirality. One of the first proposals for bringing tailored fields to chirality consisted in the use of polarization gratings that would exert an enatio-sensitive force on chiral molecules [178, 180], enabling the separation of the enantiomers (see Figure 3A, reprinted from 178). Other complex polarization spatial structures have also demonstrated to result in enantio-sensitive direction of light emission [81, 181].
@@
Let us now move towards all-optical techniques using tailored light. First, low-order non-linear interactions have the advantage of requiring moderate laser intensities. These include second-order optical techniques that break the symmetry in polarization space such as sum-frequency generation or other wave mixing effects [67, 68, 185, 186] (including the microwave regime that has been particularly successful owing to good phase matching [70, 71, 186], as shown in Figure 3C reprinted from 70) or using three resonant pulses that break the mirror symmetry in polarization and time, such as in enantio-specific state transfer measurements [179, 187] (see Figure 3D reprinted from 179) or the enantiosensitive free-induction decay steered by a tricolor cross-polarized field [84].
@@
In the highly non-linear regime, the above-mentioned chiral HHG [33] motivated additional work using tailored light to drive HHG — it was known at the time that combinations of circular drivings allow conditions of nonzero HHG, e.g. if bi-chromatic bi-circular pulses are employed [48]. The hope was to use such poly-chromatic polarization-tailored pulses to repeat this experiment but in conditions where much larger dichroism would be obtained [188]. Indeed, recently dichroism of up to  10% was measured in bicircular HHG [79, 189] (see Figure 3E reprinted from 79), and also utilized to track chiral chemical reactions in real time [190]. In the next stage, perturbative and wave-mixing methods were extended to HHG, where by tailoring the driving beams with polychromatic components the nature of the polarization-induced symmetry breaking can be more finely controlled [78]. Here one can also generate electric-dipole chiral signals that arise in the yield and handedness of the harmonics. Lastly, specialized beams where light’s symmetries are completely broken were developed, denoted as so-called locally-chiral light [53]. The idea here is to impose strict no-mirror symmetry groups on the light field at every local point in space [191, 192]. That is, the time-dependent polarization of the electric (and magnetic) field in every given point in space traces out a spatio-temporal object that cannot be superposed onto its mirror twin. This extends the typical molecular definition of chirality to electromagnetism. A direct result of this polarization-induced symmetry breaking is chiral signals in HHG (and perturbative harmonics as well) that are driven by electric-dipole interactions and arise in the yield of harmonics, rather than in their handedness (see Figure 3F, reprinted from 168). This technique has been proposed in multiple geometries, each with its own advantages/disadvantages [81, 193, 185, 82, 168, 84], and allows all-optical measurements that can be optimized to very high chiral signals, above 100%, and potentially also resolve local stereo-centers within the chiral molecule [194]. Locally-chiral light beams were also proposed as useful for extending PECD methodologies, whereby the geometry of the measurement apparatus is no longer essential to obtain the electric-dipole chiral signal since the light carries the dipole-induced symmetry breaking. This leads to a breaking of the forward/backward asymmetry in PECD and with it chiral signals in the angle-integrated above-threshold ionization spectra that can also be employed for optical enantio-purification [195].
@@
More recently, topological light beams carrying OAM were employed to develop new types of dichroism such as helical dichroism (HD) (see Figure 3G reprinted from 76). Generally, this signal appears in the linear optical response as a result of the interference between the electric dipole interaction with magnetic dipole and/or electric quadrupole interactions [196] and is a dichroism obtained upon inverting the beam’s OAM handedness (i.e. the sign of its topological charge) rather than its polarization handedness [74, 75, 76, 77]. Helical dichroism can even appear with linearly polarized beams, decoupling the need for polarization control [197]. Its origin of symmetry breaking is still in light’s spatial structure, but embedded in the phase front rather than polarization evolution. Technically, CD and HD can be combined to provide additional chiral information by controlling both degrees of freedom. In principle, nonlinear responses should also yield to HD (due to the symmetry-breaking nature being the same), which, to our knowledge, has not been generally shown or predicted yet, but only analyzed in nonlinear absorption spectroscopy [198]. Moreover, combining OAM beams with locally-chiral light can produce even more powerful chiral spectroscopy schemes [199].
In addition, recent studies show how light with evermore intricate waveforms such as three-color pulses could contribute towards an enantio-selective photochemistry via enantio-sensitive population transfers [200]. Finally, it is worth mentioning that locally-chiral light can also be used to imprint chirality in atoms [85], which are naturally achiral, by creating a superposition of electronic states that results in a chiral state [201] (see Figure 3H reprinted from 85).